A math class has 29 students. There are 3 more boys than girls. How many boys and girls are there?
Answer by Nate(3500)

plug: g = b - 3
into
g + b = 29
looks like:
b - 3 + b = 29
2b = 32
b = 16
g = b - 3 = 16 - 3 = 13